Consider the following formula:
a = 1/2 b - 4

which of the following statements is true for this formula?
A. When the value of b is less than 8, a is negative.
B. When the value of b is greater then 8, a is negative.
C. When the value of b is less than 8, a is positive.
D. When the value of b is greater then 4, a is positive.

Answer by edjones(8007)

a = 1/2 b - 4
.
A)
Let b=4
4/2-4=-2
Let b=8
8/2-4=0
True
.
Ed